# ProFL: Performative Robust Optimal Federated Learning

### v1

@@ Abstract
Performative prediction (PP) is a framework that captures distribution shifts that occur during the training of machine learning models due to their deployment. As the trained model is used, its generated data could cause the model to evolve, leading to deviations from the original data distribution. The impact of such model-induced distribution shifts in the federated learning (FL) setup remains unexplored, despite being increasingly likely to transpire in real-life use cases. Although Jin et al., (2024) recently extended PP to FL in a straightforward manner, the resulting model only converges to a performative stable point, which may be far from optimal. The methods in Izzo et al., (2021); Miller et al., (2021) can find a performative optimal point in centralized settings, but they require the performative risk to be convex and the training data to be noiseless—assumptions often violated in realistic FL systems. This paper overcomes all of these shortcomings and proposes Performative robust optimal Federated Learning (ProFL), an algorithm that finds performative optimal points in FL from noisy and contaminated data. We present the convergence analysis under the Polyak-Lojasiewicz condition, which applies to non-convex objectives. Extensive experiments on multiple datasets validate our proposed algorithms’ efficiency.
Keywords Federated learning  $\cdot$ Performative prediction $\cdot$ Robust
@@ 1 Introduction
Federated learning (FL) is a distributed learning paradigm that enables efficient learning from heterogeneous clients. Conventional FL (e.g., collaborative model training) often relies on the unrealistic assumption that each client’s data distribution remains static throughout the training process or between the training and testing phases. Recent work addresses this weakness by training models on time-evolving data (Guo et al.,, 2023; Ma et al.,, 2022) or training models on static data but deploying them on different data distributions (Nguyen et al.,, 2022; Jiang and Lin,, 2022). However, these works assume that the changes in data distribution are exogenous (i.e., are independent of the FL system). Examples of such exogenous distribution shifts include variations in data quality, lighting conditions, dynamic climate data, etc.
In many real-world applications, data often undergoes performative endogenous shifts induced by the model’s deployment. For instance, content recommended by machine learning algorithms on digital platforms can reshape consumer preferences, steering them toward modes of consumption that are easier to predict and monetize (Dean and Morgenstern,, 2022). Similarly, navigation systems can influence traffic patterns, while individuals may adapt their behavior to game decision-making systems, such as those used for loan approvals or job applications (Hardt et al.,, 2016; Zhang et al.,, 2022). These types of shifts have been studied in centralized settings under the framework of performative prediction (PP), which seeks performative stable (Perdomo et al.,, 2020) or optimal (Izzo et al.,, 2021; Miller et al.,, 2021) solutions. Addressing endogenous shifts in federated learning (FL) is not a simple extension of these centralized efforts and remains a significant challenge. Failing to do so can lead to suboptimal performance or even non-convergence.
@@
Another challenge in practical FL systems, beyond model-induced endogenous shifts, is data contamination. Client data may be noisy, corrupted, or maliciously manipulated. Unlike centralized systems, in FL, such data cannot be removed before training, as it remains with the clients. Developing FL systems that are robust to endogenous shifts and data contamination remains a significant challenge.
The only work to date that addresses the endogenous shifts in FL is Jin et al., (2024). Their approach assumes noiseless client data and only finds a performative stable point, which may be suboptimal. While Izzo et al., (2021); Miller et al., (2021) achieve a performative optimal point, their methods are centralized, assume noiseless data, and require convex performative risk. As we will show, the method that directly extends Izzo et al., (2021) to federated settings is highly vulnerable to data contamination and performs poorly when distribution shifts are complex.
@@
In this paper, we solve all these challenges simultaneously and propose Performative robust optimal Federated Learning (ProFL), an FL algorithm that finds a performative optimal solution under endogenous distribution shifts with contaminated data. We provide a theoretical analysis of data contamination and the convergence of ProFL. Compared to the federated extension of Izzo et al., (2021), ProFL employs simple yet effective strategies and converges to the performative optimal solution for a larger class of (non-convex) objective functions. In Appendix B, we provide more related work and highlight differences.
Summary of contributions. Our key contributions compared to prior works are:
  1. 1.
@@
    Finding the optimal solution. Unlike Jin et al., (2024), which only guarantees convergence to the stable point, ProFL finds the optimal point.
  2. 2.
@@
    Convergence analysis for a larger class of objectives. Unlike prior works (Izzo et al.,, 2021; Jin et al.,, 2024) that show the convergence of the proposed algorithms for the convex performative risk, we show the convergence of ProFL under Polyak-Lojasiewicz (PL) condition, which is non-convex.
  3. 3.
    Robustness to data contamination. We analyze the impact of contaminated data on the FL system and show that ProFL effectively tackles complex distribution shifts and is more resilient to estimation error and data contamination.
2 Problem Formulation
@@
Consider a set of clients $\mathcal{V}=\{1,\ldots,N\}$ collaboratively training an FL model $\theta\in\Theta\subset\mathbb{R}^{d}$ without sharing their local data. Here, $\Theta$ is a compact set (closed and bounded), and $d$ represents the model’s dimension. Let $\mathcal{Z}$ denote the sample space. Throughout training [^1] One can also interpret training as retraining or fine-tuning of models after deployment., each client interacts with the local model, causing its data to evolve based on the perceived model. Such model-induced data distributions, studied in PP (Perdomo et al.,, 2020), are characterized by a performative distribution map $D_{i}:\Theta\to\mathcal{P}(\mathcal{Z})$. Specifically, deploying model $\theta$ at client $i$ results in data following distribution $D_{i}(\theta)$. The goal is to train a global model that minimizes the total loss across all clients:
@@
where $\mathcal{L}_{i}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\ell(Z_{i};\theta)]$ is the local performative risk of client $i$ for a given loss function $\ell(Z_{i};\theta)$ on local data $Z_{i}\sim D_{i}(\theta)$. This is called performative risk because, instead of measuring the risk over a fixed distribution, it measures the expected loss of model $\theta$ on the data distribution $D_{i}(\theta)$ it induces. $\mathcal{L}(\theta)$ is the performative risk (PR) of all clients. Here, $\alpha_{i}\geq 0$ represents the fraction of client $i$’s data relative to the total, with $\sum_{i\in\mathcal{V}}\alpha_{i}=1$. The solution ${\theta}^{\text{PO}}$ that minimizes performative risk is called performative optimal (PO) point (Perdomo et al.,, 2020).
In realistic FL systems, each client’s data distribution $D_{i}(\theta)$ is unknown, and the training data may be noisy, corrupted, or maliciously manipulated. To account for this, we assume each client $i$ has a fraction $\epsilon_{i}\in[0,1]$ of data samples drawn from another fixed but arbitrary distribution $Q_{i}$. Thus, the data acquired from client $i$ follows a mixture distribution ${Z}_{i}\sim P_{i}(\theta)=(1-\epsilon_{i})D_{i}(\theta)+\epsilon_{i}Q_{i}$. We term $Q_{i}$ data “contamination”.
Our goal is to (i) design FL algorithms that converge to the performative optimal ${\theta}^{\text{PO}}$; (ii) study the impact of data contamination $Q_{i}$ and develop robust solutions. While data contamination has been studied in FL with static data, its effect in performative settings is unexplored. A moment’s thought reveals that when client data depends on the FL system, contamination in one client may cause a cascading effect across the network, making it crucial to investigate how this disrupts FL convergence. We will explore this in the paper. We next present the metrics in PP.
@@
Performative stable point. Since the data distribution $D_{i}(\theta)$ depends on $\theta$, which is the variable being optimized, finding the performative optimal point is often challenging. Thus, existing works (Perdomo et al.,, 2020; Mendler-Dünner et al.,, 2020; Zhao,, 2022) have mostly focused on finding the so-called performative stable (PS) point. In an FL system, the PS point ${{\theta}}^{\text{PS}}$ is defined as follows:
${{\theta}}^{\text{PS}}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}({{\theta}}^{\text{PS}};\theta):=\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}({{\theta}}^{\text{PS}};\theta)$ (2)
where $\mathcal{L}_{i}({{\theta}}^{{}^{\prime}};\theta)=\mathbb{E}_{Z_{i}\sim D_{i}({{\theta}}^{{}^{\prime}})}\left[\ell(Z_{i};\theta)\right]$ is called the decoupled performative risk (DPR) of client $i$. Unlike performative risk (PR), where the distribution $D_{i}(\theta)$ depends on $\theta$, the DPR decouples the two; i.e., the data distribution is induced by ${\theta}^{\text{PS}}$ while the variable being optimized is $\theta$. Since ${\theta}^{\text{PS}}$ is the fixed point of Eq. (2), it can be found via repeated risk minimization (RRM) or repeated gradient descent (RGD) (Perdomo et al.,, 2020). RRM recursively finds a model $\theta^{t+1}$ that minimizes DPR on the distribution $D(\theta^{t})$ induced by the previous model, whereas RGD recursively finds a model $\theta^{t+1}$ through gradient descent, and $\theta^{t+1}$ is not necessarily the minimizer of the DPR on $D(\theta^{t})$.
Following the idea of RGD, Jin et al., (2024) proposed the performative federated learning (PFL) algorithm to find the PS point ${\theta}^{\text{PS}}$ in FL. In PFL, each client $i$ updates its local model to reduce local DPR based on the previous model’s induced distribution: $\theta_{i}^{t+1}\leftarrow\theta_{i}^{t}-\eta_{t}\nabla\ell(Z_{i}^{t+1};\theta_{i}^{t})$, where $Z_{i}^{t+1}\sim D_{i}(\theta_{i}^{t})$ is drawn from the uncontaminated distribution. These updates are repeated for several epochs before being sent to the server for global aggregation. Jin et al., (2024) showed that PFL converges to ${\theta}^{\text{PS}}$ under clean data. Although the PS point is easier to find than the PO point, prior works (Izzo et al.,, 2021; Miller et al.,, 2021) show that the PS point may be far from optimal or may not even exist. Thus, algorithms like PFL, which target the PS point, are not ideal.
[…]

### v2

@@ Abstract
Performative prediction is a framework that captures distribution shifts that occur during the training of machine learning models due to their deployment. As the trained model is used, data generation causes the model to evolve, leading to deviations from the original data distribution. The impact of such model-induced distribution shifts in federated learning is increasingly likely to transpire in real-life use cases. A recently proposed approach extends performative prediction to federated learning with the resulting model converging to a performative stable point, which may be far from the performative optimal point. Earlier research in centralized settings has shown that the performative optimal point can be achieved under model-induced distribution shifts, but these approaches require the performative risk to be convex and the training data to be noiseless—assumptions often violated in realistic federated learning systems. This paper overcomes all of these shortcomings and proposes Performative Robust Optimal Federated Learning, an algorithm that finds performative optimal points in federated learning from noisy and contaminated data. We present the convergence analysis under the Polyak-Lojasiewicz condition, which applies to non-convex objectives. Extensive experiments on multiple datasets demonstrate the advantage of Robust Optimal Federated Learning over the state-of-the-art.
Keywords Performative prediction, Federated learning, Robust optimization, Distribution shift
@@ 1 Introduction
Federated learning (FL) is a distributed learning paradigm that enables efficient training across decentralized clients while preserving data privacy. This privacy preservation is achieved by keeping the raw training data on each client and only exchanging model updates with the central server. Such a framework has attracted significant attention in privacy-sensitive applications such as the Internet of Things, healthcare, and financial services.
Conventional FL (e.g., collaborative model training via FedAvg) typically assumes that each client’s data distribution remains static throughout the training process or between the training and deployment phases. However, this assumption often fails in real-world applications, where user behavior and data characteristics are continuously evolving. Although some previous work considers “dynamics” in FL, they largely focus on changes across clients (i.e., statistical heterogeneity) or time-varying participation rates rizk2020dynamic; park2021tackling; wang2022unified; zhu2021diurnal, rather than on distributional shifts within each client’s local data over time.
Data shifts are common in FL deployment. For instance, periodic deployment and retraining often lead to a mismatch between the data used for training and the data encountered in the future. Recent work begins to address such settings by training on time-evolving data guo2021towards; ma2022continual, or by training on one distribution and deploying on another nguyen2022fedsr; jiang2022test. However, these works primarily consider exogenous shifts—distribution changes that occur independently of the model, such as variations in sensor quality, lighting, or environmental conditions.
In contrast, many practical FL applications involve endogenous distribution shifts, where the deployed model itself induces changes in the data distribution. For example, recommendation systems shape user preferences dean2022, navigation apps influence traffic flow, and large language models affect language usageanderson2025modelmisalignmentlanguagechange. Users may also adapt their behavior to exploit automated decision systems, such as those in hiring or credit scoring Hardt2016a; zhang2022. These model-driven shifts fall under the framework of performative prediction (PP), which seeks performative stable perdomo2020performative or optimal pmlr-v139-izzo21a; miller2021outside solutions.
@@
Extending performative prediction to the federated setting is highly nontrivial. While centralized approaches can rely on access to a global dataset and full observability of distribution shifts, FL systems must contend with decentralized and partially observable dynamics. Moreover, the interaction between distributed optimization and performative feedback can lead to suboptimal convergence or even divergence if not properly addressed. Therefore, a principled treatment of performative shifts in FL is both necessary and fundamentally different from centralized solutions. Another challenge in practical FL systems, beyond model-induced endogenous shifts, is data contamination. Client data may be noisy, corrupted, or maliciously manipulated. Unlike centralized systems, in FL, such data cannot be removed before training, as it remains with the clients. Developing FL systems that are simultaneously robust to endogenous shifts and data contamination remains a significant challenge.
1.1 Related Work
Performative prediction perdomo2020performative was first introduced in 2020 as an optimization framework to deal with endogenous data distribution shifts, where the model deployed to the environment affected the subsequent data, resulting in the collected data changing as the deployed model changed. Common applications of performative prediction included strategic classification Hardt2016a; xie2024nonlinear and predictive policing ensign2018runaway. Early work focused primarily on finding performative stable (PS) points. For example, perdomo2020performative proposed Repeated Risk Minimization to find the PS point and bounded the distance between the PS point and the performative optimal (PO) point. mendler2022anticipating designed the first algorithm to find the performative stable point under the online learning setting.mendler2020stochastic proved convergence of greedy and lazy deployment under smoothness and strong convexity.
Subsequent work moved toward directly identifying PO point. miller2021outside optimized performative risk directly, although the scope of the distribution map was restricted, and pmlr-v139-izzo21a designed an algorithm, performative gradient descent, for finding the performative optimal solution. Zhao2022OptimizingTP relaxed the assumption of strong convexity to the weakly-strongly convex case and proved the convergence of the deployment. somerstep2024learning; bracale2024learning focused on learning the distribution map in performative prediction and proposed a framework named reverse causal PP. In addition, PP was also related to reinforcement learning pmlr-v162-zheng22c and bandit problems chen2023performative.
To the best of our knowledge, jin2023performative is the only work discussing performativity under the distributed setting, where they generalized Fedavg to Performative Federated Learning (PFL) and proved the uniqueness of the performative stable solution found by the algorithm. They also quantified the distance between the performative stable solution and the performative optimal solution. While the work is a breakthrough as the first to take model-dependent distribution shifts into account under the FL setting, it is only a strict generalization of perdomo2020performative without the ability to find the performative optimal point and deal with the contaminated data.
1.2 Contributions
@@
In this paper, we solve all these challenges simultaneously and propose Performative robust optimal Federated Learning (ProFL), an FL algorithm that finds a performative optimal solution under endogenous distribution shifts with contaminated data. We provide a theoretical analysis of data contamination and the convergence of ProFL. Compared to the federated extension of pmlr-v139-izzo21a, ProFL employs simple yet effective strategies and converges to the performative optimal solution for a larger class of (non-convex) objective functions. Our key contributions compared to prior works are:
@@
  • Performative the optimal solution. Unlike jin2023performative, which only guarantees convergence to the stable point, our ProFL finds the optimal point.
@@
  • Convergence analysis for a larger class of objectives. Unlike prior works pmlr-v139-izzo21a; jin2023performative that show the convergence of the proposed algorithms for the convex performative risk, we show the convergence of our ProFL converges under non-convex Polyak-Lojasiewicz (PL) condition.
  • Robustness to distribution shifts and data contamination. Our ProFL algorithm is designed for realistic FL settings with heterogeneous clients and limited observability. It integrates a principled outlier detection mechanism to enhance Robustness against client-side contamination and estimation errors. Despite operating without access to raw data, ProFL consistently achieves reliable convergence under dynamic and heterogeneous FL environments with contaminated data.
2 System Model
2.1 Performative Prediction Model
2.1.1 Centralized Performative Prediction Model
UserData $D(\theta^{t})$Model $\theta^{t+1}$PredictionsPredictions affect behavior, which shifts data distribution
Figure 1: System model of performative prediction. The model $\theta$ influences the data distribution $D(\theta)$, forming a feedback loop that drives endogenous data shifts.
Performative prediction addresses settings where the deployed model influences future data distributions (see Fig. 1). Let $\theta\in\Theta\subseteq\mathbb{R}^{d}$ denote the model parameters, where $\Theta$ is closed and convex and $d$ is the dimension. The data distribution is $\mathcal{D}(\theta)$ which depends on $\theta$ and a distribution map $\mathcal{D}(\cdot)$. $z\in\mathcal{Z}$ denotes a sample from the sample space $\mathcal{Z}$. Denote the loss function with model $\theta$ and data sample $z$ as $\ell(z;\theta)$. PP considers two risk formulations. The performative risk (PR)is defined as
$PR(\theta):=\mathbb{E}_{z\sim\mathcal{D}(\theta)}[\ell(z;\theta)].$
This is called performative risk because, instead of measuring the risk over a fixed distribution, it measures the expected loss of model $\theta$ on the data distribution $D(\theta)$ it induces. While the decoupled performative risk (DPR) is given by
$DPR(\theta,\theta^{\prime}):=\mathbb{E}_{z\sim\mathcal{D}(\theta)}[\ell(z;\theta^{\prime})].$
Unlike PR, where $D_{i}(\theta)$ depends on $\theta$, the DPR decouples the two. Based on the definitions of DPR and PR, we define two key concepts: the PS and the PO points. The PS point $\theta^{\text{PS}}$ is a fixed point that captures an equilibrium state and is defined as:
$\theta^{\text{PS}}:=\arg\min_{\theta}\mathbb{E}_{z\sim\mathcal{D}(\theta^{\text{PS}})}[\ell(z;\theta)].$ (1)
The data distribution is induced by ${\theta}^{\text{PS}}$, while $\theta$ is optimized. As ${\theta}^{\text{PS}}$ is the fixed point of Eq. (1), it can be found by repeated risk minimization (RRM) or repeated gradient descent (RGD) perdomo2020performative. RRM recursively finds a model $\theta^{t+1}$ that minimizes DPR on the distribution $D(\theta^{t})$ induced by the previous model, whereas RGD recursively finds a model $\theta^{t+1}$ through gradient descent, and $\theta^{t+1}$ is not necessarily the minimizer of the DPR on $D(\theta^{t})$. The performative optimal (PO) point, $\theta^{\text{PO}}$, minimizes the performative risk (PR) and is defined as:
$\theta^{\text{PO}}:=\arg\min_{\theta}\mathbb{E}_{z\sim\mathcal{D}(\theta)}[\ell(z;\theta)].$
The key to finding the PO point is to update the model using the actual gradient of PR, $\nabla\mathcal{L}(\theta)$, also known as the performative gradient (PG). Assume the performative distribution $D(\theta)$ has a density function $p(z;f(\theta))$, where the parameter $f(\theta)$ is a differentiable function of $\theta$. Then, the performative risk can be written as
$\mathcal{L}(\theta)=\int\ell(z;\theta)\,p(z;f(\theta))\,dz,$
provided that $p$ and $f$ are continuously differentiable. [^1] An example of $D(\theta)$ is a mixture of Gaussians, $\sum_{k=1}^{K}\nu_{k}\mathcal{N}(f_{k}(\theta),\sigma_{k}^{2})$, where the mean depends on the model pmlr-v139-izzo21a. This can approximate any smooth density to arbitrary precision.
We now present the formal problem setting of federated performative prediction, where the objective is to find the global performative optimal point in the presence of heterogeneous data shifts and possible data contamination.
2.1.2 Federated Performative Prediction Model
@@
Consider a set of clients $\mathcal{V}=\{1,\ldots,N\}$ collaboratively training an FL model $\theta$ without sharing their local data. Such model-induced data distributions, introduced in PP perdomo2020performative, are characterized by a performative distribution map $D_{i}:\Theta\to\mathcal{P}(\mathcal{Z})$. Specifically, deploying model $\theta$ at client $i$ results in data following distribution $D_{i}(\theta)$. The distribution map $D_{i}(\cdot)$ can vary across clients, resulting in data heterogeneity in the federated learning system. The goal is to train a global model that minimizes the total loss across all clients:
@@
where $\mathcal{L}_{i}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\ell(Z_{i};\theta)]$ is the PR of client $i$ for a given loss function $\ell(Z_{i};\theta)$ on local data $Z_{i}\sim D_{i}(\theta)$. Here, $\alpha_{i}\geq 0$ represents the fraction of client $i$’s data relative to the total, with $\sum_{i\in\mathcal{V}}\alpha_{i}=1$. In the FL setting, the PR over all clients is denoted by $\mathcal{L}(\theta)$, and the performative optimal point $\theta^{\text{PO}}$ is defined as its minimizer.
@@
Since the data distribution $D_{i}(\theta)$ depends on $\theta$, which is the variable being optimized, finding the PO point is often challenging. Thus, existing works perdomo2020performative; mendler2020stochastic; Zhao2022OptimizingTP have mostly focused on finding the PS point. In an FL system, the PS point ${{\theta}}^{\text{PS}}$ is defined as follows:
${{\theta}}^{\text{PS}}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}({{\theta}}^{\text{PS}};\theta):=\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}({{\theta}}^{\text{PS}};\theta),$ (3)
where $\mathcal{L}_{i}({\theta}^{\prime};\theta)=\mathbb{E}_{Z_{i}\sim D_{i}({\theta}^{\prime})}[\ell(Z_{i};\theta)]$ denotes client $i$’s DPR.
2.2 Data Contamination Model
In realistic FL systems, each client’s data distribution $D_{i}(\theta)$ is unknown, and the training data may be noisy, corrupted, or maliciously manipulated. In this work, we consider a general contamination model based on Huber’s $\epsilon$-contamination framework huber1992robust, which serves as a foundation for robust statistical analysis in the presence of outliers and data corruption. This model assumes that besides the data distribution we aim to learn, a small proportion of samples can come from an arbitrary distribution. This contamination framework is crucial for creating machine learning algorithms robust to outliers, thereby ensuring more reliable analysis in practical scenarios where data imperfections are common. Huber’s model has been fundamental in the field of robust statistics, influencing a wide range of applications and subsequent research. To account for this, we assume each client $i$ has a fraction $\epsilon_{i}\in[0,1]$ of data samples drawn from another fixed but arbitrary distribution $Q_{i}$. Thus, the data acquired from client $i$ follows a mixture distribution ${Z}_{i}\sim P_{i}(\theta)=(1-\epsilon_{i})D_{i}(\theta)+\epsilon_{i}Q_{i}$. We term $Q_{i}$ data “contamination”.
2.3 Research Objectives and Challenges
Our goal is to design FL algorithms that converge to the performative optimal ${\theta}^{\text{PO}}$ even with contaminated data $Z_{i}\sim P_{i}(\theta)$. While data contamination has been studied in FL with static data, its effect in performative settings is unexplored. When client data depends on the FL system, contamination in one client may cause a cascading effect across the network, making it crucial to investigate how this disrupts FL convergence. We will explore this in the paper.
Remark 2.1.
3 Algorithm Design
3.1 Motivation: From Stable to Optimal in Federated PP
Most existing works in the FL or decentralized setting perdomo2020performative; mendler2020stochastic; Zhao2022OptimizingTP focus on finding the PS point, which does not minimize the true PR. This paper aims to find the PO point—the model that minimizes the PR under its own induced distribution. This goal is significantly more difficult to achieve in FL than in centralized settings due to client heterogeneity, limited data and computational resources, and data contamination, which further distort the model-dependent distributions. To the best of our knowledge, no FL algorithm has been shown to converge to the PO point, even under clean data and convex loss assumptions. This paper fills that gap by proposing an FL algorithm that provably converges to the PO point, accommodates non-convex (PL) conditions, and is robust to data contamination.
3.2 The Proposed Algorithm
In this section, we propose several simple yet effective strategies that collectively lead to a robust federated performative prediction algorithm, called Performative robust optimal Federated Learning (ProFL). The pseudocode is given in Algorithm 1. Unlike jin2023performative that focus on PS point, ProFL aims to achieve convergence to the global PO point in federated PP. Achieving the PO point requires estimating the PG, which captures both the loss and the feedback from model-induced distribution shifts. We begin by describing how the PG is estimated in the federated setting and then introduce our strategies for adapting this process to the challenges of FL, including client heterogeneity, limited data and computational resources, and data contamination.
3.2.1 Federated Estimation of the Performative Gradient
The key to finding the PO point is to update the model using the performative gradient (PG), which captures both the direct loss gradient and the effect of model-induced distributional shifts. While pmlr-v139-izzo21a proposed a method to estimate the PG, their analysis is limited to a centralized setting with a specific distribution map $D(\theta)$, where $f(\theta)$ is assumed to be a linear function of $\theta$, and the performative risk (PR) is convex. However, such assumptions are restrictive and often fail to hold in practice, especially under data contamination. In FL, the challenge becomes even greater. Unlike the centralized setting, which computes a single performative gradient $\nabla\mathcal{L}(\theta)$ on the server, FL requires each client to update its model using a local PG: $\nabla\mathcal{L}_{i}(\theta)$. Assume the local performative distribution $D_{i}(\theta)$ has density $p(z;f_{i}(\theta))$ with parameter $f_{i}(\theta)$ being a function of $\theta$, then $\mathcal{L}_{i}(\theta)=\int\ell(z;\theta)p(z;f_{i}(\theta))dz$ when functions $p$ and $f_{i}$ are continuously differentiable [^2] An example of $D_{i}(\theta)$ is a mixture of Gaussians, $\sum^{K}_{k=1}\nu_{k}\mathcal{N}(f_{k}(\theta),\sigma_{k}^{2})$, where the mean depends on the model pmlr-v139-izzo21a. This can approximate any smooth density distribution to arbitrary precision. . PG can be derived